graph the function g(x)=-3x-1
Answer by addingup(3677)

g(x)=-3x-1
Rewrite:
y = -3x-1
Remember the slope-intercept form is y = mx+b, so here we have that your slope is -3 and your intercept is -1. To get an ordered pair we can graph:
Set x to 0:
y = -3(0)-1 = 0-1 we have a pair (0,-1)
now using algebra get the x out and the y in:
y+3x = -1
3x = -y-1
x = (-y-1)/3 set y to 0
x = -0-1/3 now we have a second pair (-1/3, 0) Note: the ordered pairs are always written x, y.